OpenStudy (anonymous):

Solve: -1/2x + 1 = −x + 8 7 14 18 1/2

OpenStudy (anonymous):

the fractions are messing me up

OpenStudy (anonymous):

-1/2x + 1 = −x + 8 multiply everything by 2 to remove the fraction -1x + 2 = -2x + 16 -1x + 2x = -2 + 16 x = 14